# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

### v3

@@ 1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable versatility and modeling capabilities. However, several challenges still hinder their widespread applicability, with scalability being the most significant concern. This scalability issue affects both the training and inference phases. Previous research has sought to mitigate the higher training computation by using three broad categories of methods: 1. Training every iteration on the sampled subgraph 2. Training on a coarsened graph generated from a coarsening algorithm 3. Training on a synthetic graph that mimics the original graph . Unfortunately, these smaller graphs can only be utilized for training purposes and have not alleviated the computational burden during inference since the method mentioned above requires the whole graph during the inference phase. Consequently, for larger graphs, not only do inference times increase, but memory consumption may also exceed limitations.
@@ 2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by Chiang et al. [1] and Zeng et al. [2] uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen and Zhu [3], Chen et al. [4], Cong et al. [5], Zou et al. [6]. Additionally, Fahrbach et al. [7] introduced a coarsening algorithm based on Schur complements to create embeddings for vertices in large graphs, while Huang et al. [8] (SGGC) applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. Kumar et al. [9] proposed feature-based coarsening algorithms. Jin et al. [10] (GCOND), Wang et al. [11] (GC-SNTK) presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. Gupta et al. [12] (BONSAI) addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. Later, Xue et al. [13] shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin et al. [14] (DOSCOND) and Xu et al. [15] (KIDD) tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
@@ 3.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. Kipf and Welling [16] introduced Graph Convolutional Network (GCN) as follows:
@@ 3.3 Graph Coarsening and Graph Partitioning
Loukas [17] discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. SGGC used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
@@
Figure 2: Methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
@@ 4 Methodology
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. Xue et al. [13] introduced this approach to reduce information loss after partitioning. We add an edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. Liu et al. [18] introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges but do not sample from the constructed set of subgraphs.
$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\quad\text{and}\quad\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (2)
@@
Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
Proof.
In Appendix A.1. ∎
@@
  • From Equation 2, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$ This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
@@ 4.3 Time and Space Complexity
The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X^{(l)}$ and $(d\times d)$ for $\mathcal{W}^{(l)}$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$. The complexity is based on GCN. For other architectures like GAT [19], the complexity changes.
@@
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph. Table 1 compares different approaches.
@@
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}.$
@@
$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]$
$=k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d$
$=nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\quad(\textup{where }k=nr)$
$=nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$. Using this, we can say,
$nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\leq nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d=nd^{2}+n^{2}d\\$
Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$. ∎
For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches. In Section C, we show that, for datasets with millions of nodes, the conditions outlined in Lemma 4.2 can indeed be satisfied. As the number of nodes increases, the gap in inference time between baseline methods and FIT-GNN widens. In table 9, we show the upper bound of $\phi_{\max}$ calculated as per the conditions mentioned in Lemma 4.2 for different datasets corresponding to different coarsening ratios. Negative upper bound is observed for some datasets, implying that our method won’t give better computational complexities in those scenarios.

### v4

@@ 1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable versatility and modeling capabilities. However, several challenges still hinder their widespread applicability, with scalability being the most significant concern. This scalability issue affects both the training and inference phases. Previous research has sought to mitigate the high training computation cost by using three broad categories of methods: 1. Training every iteration on a sampled subgraph 2. Training on a coarsened graph generated from a coarsening algorithm 3. Training on a synthetic graph that mimics the original graph . Unfortunately, these smaller graphs can only be utilized for training purposes and have not alleviated the computational burden during inference since the method mentioned above requires the whole graph during the inference phase. Consequently, for larger graphs, not only do inference times increase, but memory consumption may also exceed limitations.
@@ 2 Related Work
The aim of reducing computational costs has been approached through various methods. Research by 10.1145/3292500.3330925 and Zeng2020GraphSAINT: uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training. Other methods, including layer-wise sampling and mini-batch training, have been explored by chen2018stochastic; chen2018fastgcn; 10.1145/3394486.3403192; DBLP:journals/corr/abs-1911-07323. Additionally, DBLP:journals/corr/abs-2007-02817 introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while 10.1145/3447548.3467256 [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time. pmlr-v202-kumar23a; joly2024graph; NEURIPS2024_733209a1; 10.1145/3589335.3651920 also proposed different coarsening algorithms for scalable training approaches. jin2021graph [GCOND], wang2024fast [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph. However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden. gupta2025bonsai [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees. However, these approaches do not decrease inference time since the inference phase takes the whole graph as input. DBLP:journals/corr/abs-2202-00075 shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach. In the graph-level task, Jin_2022 [DOSCOND] and 10.1145/3580305.3599398 [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK. However, the graph regression problem was not explored in this direction.
@@ 3.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors. kipf2017semisupervised introduced Graph Convolutional Network (GCN) as follows:
@@ 3.3 Graph Coarsening and Graph Partitioning
loukas2018graphreductionspectralcut discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. $X^{\prime}$ is the coarsened node representation. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $n\times r, kk\in\mathbb{Z}$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$. $A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. SGGC (10.1145/3447548.3467256) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
@@
Figure 2: Figure showing the comparison between the Extra Node Method and the Cluster Node Method of appending additional nodes in $G_{1},G_{2},G_{3}$.
@@ 4 Methodology
  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$. DBLP:journals/corr/abs-2202-00075 introduced this approach to reduce information loss after partitioning. We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$. Node embedding of $v\in\mathcal{E}_{G_{i}}$ is $x_{v}\in X$.
    $\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$ (2)
  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist. Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$. liu2024scalable introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph. In our work, we add cross-cluster edges. For each node $t\in\mathcal{C}_{G_{i}}$, the node embedding $x_{t}=X^{\prime}_{t}$.
    $\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (3)
@@
Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used. (Proof in Appendix A.1)
@@
  • From Equation 2 and 3, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$. This is because, using Extra Nodes, we add all the neighboring nodes. However, using Cluster Nodes, we only add one representative node for each neighboring cluster. Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
@@ 4.3 Time and Space Complexity
In Equation 1, the matrix dimensions are as follows: $\tilde{A}$ has dimensions $(n\times n)$, $X^{(l)}$ is of dimensions $(n\times d)$, and $\mathcal{W}^{(l)}$ has dimensions $(d\times d)$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$. The complexity is based on GCN. For other architectures like GAT veličković2018graphattentionnetworks, the complexity changes.
@@
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{i}$ and $\phi_{i}$ is the number of additional nodes appended in each subgraph. Table 1 compares different approaches.
@@
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$ if $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}$.
@@
$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{i}^{2}d+2n_{i}\phi_{i}d+n_{i}d^{2}+\phi_{i}d^{2}]$
$=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
Now, define $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$. We also define the variances and covariance,
$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2}$ $\qquad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},$
$\text{Cov}(n_{i},\phi_{i})=\text{Cov}$ $:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
We can also write,
$\text{Var}(n_{i}+\phi_{i})=\frac{1}{k}\sum_{i=1}^{k}(n_{i}+\phi_{i})^{2}-(\mathbb{E}[n_{i}]+\mathbb{E}[\phi_{i}])^{2}=\text{Var}(n_{i})+\text{Var}(\phi_{i})+2\text{Cov}(n_{i},\phi_{i})$
Using the definitions above, we can say:
$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
$=nd^{2}+d\left[k(\textup{Var}(n_{i})+\mathbb{E}[n_{i}]^{2})+k(\textup{Var}(\phi_{i})+\mathbb{E}[\phi_{i}]^{2})+2k(\textup{Cov}(n_{i},\phi_{i})+\mathbb{E}[n_{i}]\mathbb{E}[\phi_{i}])+dk\mathbb{E}[\phi_{i}]\right]$
$=nd^{2}+nd\left[r\textup{Var}(n_{i})+\frac{1}{r}+r\textup{Var}(\phi_{i})+r\mathbb{E}[\phi_{i}]^{2}+2r\textup{Cov}(n_{i},\phi_{i})+2\mathbb{E}[\phi_{i}]+rd\mathbb{E}[\phi_{i}]\right]$
$=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
Given, $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}\Rightarrow\mathbb{E}[\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-(\frac{d}{2}+\frac{1}{r})$. Using this, we can say the following:
$\sum_{i=1}^{k}[(n_{i}+$ $\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
$\leq nd^{2}+nd\left[r\left(\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})+\frac{d^{2}}{4}+\frac{1}{r^{2}}+\frac{d}{r}-2\left(\frac{d}{2}+\frac{1}{r}\right)\sqrt{\Delta}\right)\right$.
$\qquad\qquad\left.+2\sqrt{\Delta}-2\left(\frac{d}{2}+\frac{1}{r}\right)+rd\sqrt{\Delta}-rd\left(\frac{d}{2}+\frac{1}{r}\right)+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
$\qquad\qquad\qquad(\textup{where }\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})=\Delta)$
$=nd^{2}+n^{2}d$
Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$. ∎
Corollary 4.3.
$\mathbb{E}[\phi_{i}]$ has a positive upper bound when $\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$.
Proof.
From Lemma 4.2, we know the upper bound of $\mathbb{E}[\phi_{i}]$ is $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)$.
If $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$, then we can say the following:
$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$
$\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$
$\Rightarrow\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$
For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches. From Corollary 4.3, we can understand that it is ideal to have similarly sized subgraphs. For more details on time and space complexity, refer to Section C in the Appendix.